OpenStudy (anonymous):

If a = 2i + 3j, b = 3i – 4j, and c = 2i + j find y and u such that c = ya + ub

OpenStudy (anonymous):

Help pleeeeeaaaasssseeee

OpenStudy (anonymous):

plug in a,b,c into c=ya+ub yields 2i+j= y(2i+3j) + u(3i-4j). therefore 2i+j= 2yi+3yj + 3ui -4uj. --> 2i+j = (2y+3u)i +(3y-4u)j equate both sides to have 2 equations: 1) 2=2y+3u 2) 1= 3y-4u . solve these equations to get y and u

OpenStudy (anonymous):

thankyou :D

OpenStudy (anonymous):

your welcome